OpenStudy (anonymous):

integrate (1/(sqrt(x^2+10))dx

ganeshie8 (ganeshie8):

trig substitution

ganeshie8 (ganeshie8):

maybe, sub \(x = \sqrt{10} \tan \theta\)

OpenStudy (anonymous):

for substitution, what is your u?

OpenStudy (anonymous):

@ganeshie8 ^^^^

hartnn (hartnn):

so, you're comfortable with 'u' from ganeshie's sub., x = sqrt 10 tan u gives \( u = \tan^{-1}(x/\sqrt {10}) \)

ganeshie8 (ganeshie8):

\(\large \mathbb{ \int \frac{1}{\sqrt{x^2+10}}}dx\) sub \(x = \sqrt{10} \tan \theta\) \(dx = \sqrt{10} \sec^2 \theta d\theta\) \(\large \mathbb{ \int \frac{\sqrt{10}\sec^2\theta d\theta}{\sqrt{10 \tan^2\theta +10}}}\) \(\large \mathbb{ \int \frac{\sec^2\theta d\theta}{\sqrt{ \tan^2\theta +1}}}\) \(\large \mathbb{ \int \frac{\sec^2\theta d\theta}{\sqrt{ \sec^2\theta}}}\) \(\large \mathbb{ \int \sec \theta d\theta}\)

ganeshie8 (ganeshie8):

simplifies to that, after that u may pull the integral of sec as formula : \(\large \int \sec x = \ln|\sec x + \tan x| + C \)

ganeshie8 (ganeshie8):

and dont forget to convert the theta's in final answer...